y(2xy+1)dx-xdy=0


1
Expert's answer
2020-06-10T19:59:18-0400

"y(2xy+1)dx-xdy=0\\
y(2xy+1)-x\frac{dy}{dx}=0\\
y'-\frac{1}{x}y = 2y^2"

it is Bernoulli equation

"v=y^{-1}\\
-v'-\frac{v}{x}=2\\
v'+\frac{v}{x} = 0\\
\frac{dv}{v} = -\frac{dx}{x}\\
\ln(v) = -\ln(x)+C\\
v(x) = \frac{C}{x}"

general solution is:

"v(x) = -x+\frac{c}{x}\\
y(x) = v^{-1}(x) = \frac{x}{C-x^2}"